OpenStudy (annah.bee):

3. Summarize how density is changed by changing the mass without changing the volume, and by changing volume of a substance with out changing mass.

OpenStudy (annah.bee):

ill fan and medal please help

OpenStudy (anonymous):

Here's a mathematical expression for density: \[\rho =\frac{ m }{ V }\] We can see that if we increase the mass (while volume remains constant) that density also increases! Let's take m=10 and V=5 So we get \(\rho=10/5=2\) Now what about if we change m to 15 instead of 10? \(\rho=15/5=3\) Therefore we say that the density is direction proportional to the mass and inversely proportional to volume

OpenStudy (annah.bee):

thank you